OpenStudy (anonymous):

with expression is equivalent to \[\sqrt{a^2b^1/2}^{-1}\]

jhonyy9 (jhonyy9):

can you rewriting squaroot using exponent ?

jhonyy9 (jhonyy9):

so than sqrtx = x^(1/2) can using this ?

OpenStudy (anonymous):

yea

jhonyy9 (jhonyy9):

b has exponent of 1/2 inside radical ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

and -1 is the exponent of radical ?

OpenStudy (anonymous):

yeah

jhonyy9 (jhonyy9):

hope you know that exponent of -1 what mean ? for example a^(-1) = 1/a

OpenStudy (anonymous):

yea is it